. A cyclist traveled at a rate of 32 mph to visit a nearby town. The cyclist averaged 10 mph on the return trip. If the round trip took 6.3 hours, find the distance to the nearby town.
Answer by edjones(7311)

D=rt
Let t=the time it take to go to the town.
32t=10(6.3-t) since the distance is equal.
32t=63-10t
42t=63
2t=3
t=1.5 hrs
1.5*32=48 mi distance to town.
.
Ed